OpenStudy (daenio):

The Ontario Lottery and Gaming Corporation states that probability of winning $5 on a $2 scratch ticket is 1 in 10. If a person buys one scratch ticket every day for a week, the probability that he or she will win $5 twice is Question 2 options: 18% 10% 1% 12%

OpenStudy (anonymous):

\[\dbinom{7}{2} (.1)^2(.9)^5\]

OpenStudy (anonymous):

\[\dbinom{7}{2}=\frac{7\times 6}{2}=7\times 3=21\] but you need a calculator for the rest

OpenStudy (anonymous):

i get http://www.wolframalpha.com/input/?i=21*.1^2*.9^5 about 12%

OpenStudy (daenio):

Where did you get 0.9 from?

OpenStudy (daenio):

@satellite73

OpenStudy (daenio):

Oh, was the 0.9, the probability that one would lose?

OpenStudy (anonymous):

probability you win is .1, probability you lose is 1-.1=.9

OpenStudy (daenio):

Okay, I get it now! Thanks

OpenStudy (anonymous):

win twice, lose 5 times

OpenStudy (anonymous):

yw